# Non-Bloch band theory for time-modulated discrete mechanical systems

### v1

@@ Abstract
This study establishes a non-Bloch band theory for time-modulated discrete mechanical systems. We consider simple mass-spring chains whose stiffness is periodically modulated in time. Using the temporal Floquet theory, the system is characterized by linear algebraic equations in terms of Fourier coefficients. This allows us to employ a standard linear eigenvalue analysis. Unlike non-modulated linear systems, the time modulation makes the coefficient matrix non-Hermitian, which gives rise to, for example, parametric resonance, non-reciprocal wave transmission, and non-Hermitian skin effects. In particular, we study finite-length chains consisting of spatially periodic mass-spring units and show that the standard Bloch band theory is not valid for estimating their eigenvalue distribution. To remedy this, we propose a non-Bloch band theory based on a generalized Brillouin zone. The proposed theory is verified by some numerical experiments.
@@
Non-Bloch band theory , Time-modulated system , Non-reciprocal wave propagation , Floquet theory , Non-Hermitian skin effect , Non-Hermitian physics
@@ 1 Introduction
It is well known that waves propagating in linear and time-independent media obey reciprocity, i.e., transmission between any two points in space is independent of the propagation direction. In periodic lattices, this implies that Bloch bands are symmetric with respect to the inversion of the Bloch wavevector. Reciprocity holds in a number of classical and quantum systems, e.g., photonics, acoustics, and solid mechanics. Recent studies have revealed that reciprocity breaks down in nonlinear or time-varying media [1, 2, 3]. Broken reciprocity realizes, for example, asymmetric mode conversion [4, 5], unidirectional wave propagation [6, 7, 8], and topologically-protected one-way edge modes [9].
@@
To realize non-reciprocity in mechanical systems, the simplest model would be an array of masses and springs whose constants are modulated in time [10, 11, 12, 9, 13, 14, 15]. Without using any nonlinear elements, the time modulation allows us to break time-reversal symmetry and reciprocity of the system. Such time-modulated systems are theoretically tractable with the help of the temporal Floquet theory [16, 17, 18] and also experimentally feasible using magnets or gyroscopic coupling [9, 12, 7].
@@
When a system is periodic both in space and time, the Bloch–Floquet theory formulates a linear or quadratic eigenvalue problem in terms of angular (quasi)frequency $\omega$ for each Bloch wavenumber (wavevector) $k$. When the eigenvalue problem is written in the form $\mathcal{H}(k)\varphi=\omega\varphi$, the square matrix $\mathcal{H}$ (Bloch Hamiltonian) can be non-Hermitian unlike usual time-independent systems. This means that eigenvalues $\omega$ are complex and eigenvectors (eigenmodes) $\varphi$ generally do not form an orthogonal system. Non-Hermiticity gives rise to unique concepts such as parity-time (PT) symmetry [19, 20], exceptional points [21, 22, 23], and non-Hermitian skin effects, which have been intensively studied in the name of non-Hermitian physics [24].
@@
The Bloch band theory with spatial periodicity is a useful tool for estimating the spectral properties of finite-size systems. In usual Hermitian systems, it is well known that Bloch bands represent the spectrum of a lattice (chain) whose size (length) is sufficiently large. The non-Hermiticity defies this conventional knowledge. Due to the presence of non-Hermitian skin effects [25], the standard Bloch (quasi-periodic) condition fails to describe some bulk states. This causes significant differences between the spectrum of finite-size systems with open boundary conditions (OBCs) and purely periodic lattices without boundaries, which can be analyzed from the spectral topology in the complex frequency (energy) plane [26, 27].
@@
The breakdown of the Bloch band theory motivates us to establish a novel framework that is valid for general non-Hermitian systems. A promising approach is the non-Bloch band theory [28, 29] with a generalized Brillouin zone (GBZ) [25, 30]. The non-Bloch band theory allows the wavenumbers (wavevectors) to be complex such that quasi-periodic eigenmodes represent exponentially growing or decaying bulk states. The non-Bloch band theory is validated in some theoretical models (e.g. Hatano–Nelson model [31] and Su-Schrieffer-Heeger model [32]) in quantum physics.
@@
As non-Hermitian mechanical systems are also subjects of intensive study, some theories have been developed and validated. Melkani and Paulose studied time-modulated discrete mechanical systems based on space-time symmetry [33]. Ghatak et al. discussed the non-Hermitian topology in an active mechanical metamaterial in analogy to the Su-Schrieffer-Heeger model [34]. A mathematical theory for time-modulated discrete systems is also established by Ammari and Kosche [18]. Nevertheless, it remains unclear how the spectrum of time-modulated dynamic mechanical systems with OBCs is associated with that of periodic lattices.
@@
In this study, we propose a non-Bloch band theory for time-modulated discrete mechanical systems. Time-modulated mass-spring chains are formulated using the temporal Floquet theory. We present numerical evidence that the conventional Bloch band theory is not valid for the time-modulated systems. Following [29], we define a generalized Brillouin zone and establish a non-Bloch band theory. This allows us to estimate eigenvalues of long chains in the complex frequency plane even in the presence of non-Hermitian skin modes that arise from the non-reciprocity.
@@ 2.1 Model
with constants $G_{j}>0$, $\gamma_{j}\in[0,G_{j})$, and $\phi_{j}\in[-\pi,\pi]$. Since all the coefficients $\mu_{j}$ are $T$-periodic, i.e., $\mu_{j}(t+T)=\mu_{j}(t)$ with $T=2\pi/\Omega$ for all $j=1,\ldots,M$ and $t\in\mathbb{R}$, we use the Floquet theory [16, 17, 18] and seek solutions $u^{j}$ of the form
@@
with unknown coefficients $U^{j}_{n},V^{j}_{n}\in\mathbb{C}$ ($n\in\mathbb{Z}$) for each $j=1,\ldots,M$. The coefficients should satisfy $(\omega+n\Omega)U^{j}_{n}=\mathrm{i}V^{j}_{n}$ for each $j$ and $n$. Note that the response $u^{j}$ is not in general time-harmonic with angular frequency $\omega$ despite the harmonic excitation $f_{j}$. In view of this, the factor $\omega$ in 6 is called quasifrequency [17].
@@
$F=$ $(0,\ldots,F_{1},\ldots,0,\ldots,0,\ldots,F_{N},\ldots,0)^{T}.$ (19)
@@
is not necessarily Hermitian despite that each matrix is Hermitian. Thus, we cannot always expect that eigenvalues $\omega$ of $\mathcal{H}$ are all real. However, due to the Hermiticity of the two matrices, it can be proved that the eigenvalues $\omega$ are real or come in complex-conjugate pairs $(\omega,\bar{\omega})$ [35].
@@ 2.2 Numerical example
Figure 2 (b) and (c) show the trajectories of eigenvalues $\omega$ of $\mathcal{H}$ as the modulation amplitude $\delta$ varies from 00 to $0.99$. As the spectrum is $\Omega$-periodic in the complex $\omega$-plane [17], we plot eigenvalues only within the range of $0\leq\mathrm{Re}[\omega]\leq\Omega$. The results indicate that all the eigenvalues are purely real when the modulation amplitude $\delta$ is less than $0.843$. Above this threshold, the eigenvalues come in a conjugate pair $(\omega,\bar{\omega})$ with nonzero imaginary part. The eigenvalues with negative and positive imaginary part are associated with modes decaying and growing in time, respectively (recall that the convention $\mathrm{e}^{-\mathrm{i}\omega t}$ is used). This means that the system undergoes a phase transition from stable to unstable states. This is a typical phenomenon in PT symmetric systems and called real-to-complex eigenvalue transition, pseudo-Hermiticity breaking, or PT phase transition [36, 33].
@@
Figure 2: Verification of the Floquet theory for a time-modulated system. (a) Two mass-spring pairs. (b), (c) Trajectory of eigenvalues $\omega$ as the modulation amplitude $\delta$ varies from 00 to $0.99$. (d) displacements $u^{1}$ and $u^{2}$ as a function of time $t$ when the initial conditions $u^{1}(0)=u_{0}$ and $u^{2}(0)=\partial_{t}u^{1}(0)=\partial_{t}u^{2}(0)=0$ are given.
Figure 3: Spectrum of $\mathcal{H}$ for a long-chain of mass-spring pairs. (a) A long chain consists of spatially periodic units. Its unit comprises three mass-spring pairs. (b) Time modulation of $\mu^{1}$, $\mu^{2}$, and $\mu^{3}$. (c) Distribution of eigenvalues $\omega$ for the system (a). Some representatives are denoted by the symbols A, B, C, D, and E with their conjugate pairs. (d) Eigenmodes corresponding to the representatives.
@@ 3 Non-Bloch band theory
In the previous section, we observed that the long chain with repeating units exhibit some special spectral structures and its eigenmodes can be localized at the edges. To analyze these unique properties, we approximate the long chain as a purely periodic system. The interesting point is that the standard Bloch band theory breaks down for our time-modulated system due to the non-Hermiticity of $\mathcal{H}$. In this section, we introduce the non-Bloch band theory [28, 29] with a generalized Brillouin zone [25, 30] to study the spectrum of $\mathcal{H}$ for long chains with OBCs.
@@ 3.1 Bloch band theory with Brillouin zone
which depends on $k$. Let $\sigma(k)=\{\omega\in\mathbb{C}:\mathrm{det}(\omega I-\mathcal{H}_{k})=0\}$ be the spectrum of $\mathcal{H}_{k}$. The union $\sigma_{\mathrm{BZ}}:=\cup_{k\in\mathbb{R}}\ \sigma(k)$ is called a Bloch band. For the computation of $\sigma_{\mathrm{BZ}}$, it suffices to consider $k\in[-\pi,\pi]$ as the $2\pi$ shift does not affect the Bloch condition 24, i.e., $\mathcal{H}_{k+2\pi}=\mathcal{H}_{k}$ and $\sigma_{\mathrm{BZ}}=\cup_{k\in[-\pi,\pi]}\ \sigma(k)$. The range $[-\pi,\pi]$ is called the (first) Brillouin zone (BZ).
@@
However, this well-known fact does not always hold true for non-Hermitian systems [27]. The Bloch condition 24 with real $k$ implicitly enforces solutions not to grow or decay in space, which is not a reasonable assumption from the results in Figure 3 (d).
To see this, we present a numerical example of the Bloch band $\sigma_{\mathrm{BZ}}$ in Figure 4. As shown in Figure 4 (a) and (b), we use the same modulation frequency $\Omega$ and material parameters $m_{j}$, $G_{j}$, $\gamma_{j}$, and $\phi_{j}$ in 21, 22 and 23. Figure 4 (c) and (d) show the Bloch band diagram on the Brillouin zone $-\pi\leq k\leq\pi$. As the periodic system is also non-Hermitian, the eigenvalues of $\mathcal{H}_{k}$ are complex-valued in general. Notably, the bands are asymmetric with respect to the inversion $k\to-k$, which implies that the system is non-reciprocal. Figure 4 (e) shows the comparison between the bands $\sigma_{\mathrm{BZ}}$ and spectrum of the long chain $\sigma_{\mathrm{OBC}}$ in Figure 3 (c). From the result, we observe the obvious discrepancy between the two spectra. This implies that the Bloch band theory with the standard BZ does not work for our time-modulated system.
@@ 3.2 Non-Bloch band theory with generalized Brillouin zone
The system 28 yields a (generalized) linear eigenvalue problem in terms of $\omega$ for fixed $\beta\in\mathbb{C}$. When $\omega\in\mathbb{C}$ is given, this is a quadratic eigenvalue problem in terms of $\beta$, which can be recast into a linear eigenvalue problem [35].
For each $\omega\in\mathbb{C}$, let $\beta_{1}(\omega),\beta_{2}(\omega),\ldots,\beta_{2N}(\omega)$ be eigenvalues of 28, where $N$ is the size of the partitioned matrices in 28. The eigenvalues are sorted such that $|\beta_{1}(\omega)|\leq|\beta_{2}(\omega)|\leq\ldots\leq|\beta_{2N}(\omega)|$. A generalized Brillouin zone (GBZ) [25, 30] is defined by
@@
To see this, for a fixed $\omega\in\mathbb{C}$ and $F=0$, we consider a solution of 12 of the form
$\begin{pmatrix}U^{3j-2}\\ V^{3j-2}\end{pmatrix}=(\beta^{(1)})^{j-1}\varphi^{(1)}+(\beta^{(2)})^{j-1}\varphi^{(2)}$ (33)
for each $j$, where $\beta^{(1)}$ and $\beta^{(2)}$ are eigenvalues of 28 with $|\beta^{(1)}|\leq|\beta^{(2)}|$. The vectors $\varphi^{(1)}$ and $\varphi^{(2)}$ should be determined from OBCs. For simplicity, let us consider the clamped boundary condition at the left and right edges, i.e., $(U^{1},V^{1})=(U^{3M-2},V^{3M-2})=0$. Then, we arrive at the characteristic equation
$\mathrm{det}\begin{bmatrix}I&I\\ (\beta^{(1)})^{M-1}I&(\beta^{(2)})^{M-1}I\end{bmatrix}=0$ (34)
or equivalently $(\beta^{(1)})^{M-1}-(\beta^{(2)})^{M-1}=0$. If $|\beta^{(1)}|<|\beta^{(2)}|$, then this implies that both $\beta^{(1)}$ and $\beta^{(2)}$ tend to 00 as $M\to\infty$. This contradicts that $\beta^{(1)}$ and $\beta^{(2)}$ are eigenvalues of 28. Therefore, we need to impose the condition $|\beta^{(1)}|=|\beta^{(2)}|$, i.e., $|\beta_{N}|=|\beta_{N+1}|$.

### v2

@@ Abstract
This study establishes a non-Bloch band theory for time-modulated discrete mechanical systems. We consider simple mass-spring chains whose stiffness is periodically modulated in time. Using the temporal Floquet theory, the system is characterized by linear algebraic equations in terms of Fourier coefficients. This allows us to employ a standard linear eigenvalue analysis. Unlike non-modulated linear systems, the time modulation makes the coefficient matrix non-Hermitian, which gives rise to, for example, parametric resonance, non-reciprocal wave transmission, and non-Hermitian skin effects. In particular, we study finite-length chains consisting of spatially periodic mass-spring units and show that the standard Bloch band theory is not valid for estimating their eigenvalue distribution. To remedy this, we propose a non-Bloch band theory based on a generalized Brillouin zone. This novel approach, the combination of the temporal Floquet theory for time-modulated systems and generalized Brillouin zone, enables the prediction of eigenvalue distrubution under open boundary conditions and also quantitative characterization of non-Hermitian skin modes. The proposed theory is verified by some numerical experiments.
@@
Non-Bloch band theory , Generalized Brillouin zone , Time-modulated system , Non-reciprocal wave propagation , Floquet theory , Non-Hermitian skin effect , Non-Hermitian physics
@@ 1 Introduction
It is well known that waves propagating in linear and time-independent media obey reciprocity, i.e., transmission between any two points in space is independent of the propagation direction. In periodic lattices, this implies that Bloch bands are symmetric with respect to the inversion of the Bloch wavevector. Reciprocity holds in a number of classical and quantum systems, e.g., photonics, acoustics, and solid mechanics. Recent studies have revealed that reciprocity breaks down in nonlinear or time-varying media sounas2017non-reciprocal , fleury2015nonreciprocal , nassar2020nonreciprocity . Broken reciprocity realizes, for example, asymmetric mode conversion zanjani2014one-way , chen2021efficient , unidirectional wave propagation liang2009acoustic , chen2019nonreciprocal , yi2018one-way , and topologically-protected one-way edge modes wang2015topological .
@@
To realize non-reciprocity in mechanical systems, the simplest model would be an array of masses and springs whose constants are modulated in time nassar2017non-reciprocal , swinteck2015bulk , wang2018observation , wang2015topological , vila2017bloch-based , melkani2023space-time , pu2024multiple . Without using any nonlinear elements, the time modulation allows us to break time-reversal symmetry and reciprocity of the system. Such time-modulated systems are theoretically tractable with the help of the temporal Floquet theory ammari2022nonreciprocal , ammari2023transmission , ammari2024topological and also experimentally feasible using magnets or gyroscopic coupling wang2015topological , wang2018observation , chen2019nonreciprocal .
@@
When a system is periodic both in space and time, the Bloch–Floquet theory formulates a linear or quadratic eigenvalue problem in terms of angular (quasi)frequency $\omega$ for each Bloch wavenumber (wavevector) $k$. When the eigenvalue problem is written in the form $\mathcal{H}(k)\varphi=\omega\varphi$, the square matrix $\mathcal{H}$ (Bloch Hamiltonian) can be non-Hermitian unlike usual time-independent systems. This means that eigenvalues $\omega$ are complex and eigenvectors (eigenmodes) $\varphi$ generally do not form an orthogonal system. Non-Hermiticity gives rise to unique concepts such as parity-time (PT) symmetry ruter2010observation , matsushima2022unidirectional , exceptional points miri2019exceptional , matsushima2023exceptional , even2024experimental , and non-Hermitian skin effects, which have been intensively studied in the name of non-Hermitian physics ashida2020non-hermitian .
@@
The Bloch band theory with spatial periodicity is a useful tool for estimating the spectral properties of finite-size systems. In usual Hermitian systems, it is well known that Bloch bands represent the spectrum of a lattice (chain) whose size (length) is sufficiently large. The non-Hermiticity defies this conventional knowledge. Due to the presence of non-Hermitian skin effects yao2018edge , the standard Bloch (quasi-periodic) condition fails to describe some bulk states. This causes significant differences between the spectrum of finite-size systems with open boundary conditions (OBCs) and purely periodic lattices without boundaries, which can be analyzed from the spectral topology in the complex frequency (energy) plane zhang2020correspondence , zhang2022review .
@@
The breakdown of the Bloch band theory motivates us to establish a novel framework that is valid for general non-Hermitian systems. A promising approach is the non-Bloch band theory kunst2019non-hermitian , yokomizo2019non-bloch with a generalized Brillouin zone (GBZ) yao2018edge , yang2020non-hermitian . The non-Bloch band theory allows the wavenumbers (wavevectors) to be complex such that quasi-periodic eigenmodes represent exponentially growing or decaying bulk states. The non-Bloch band theory is validated in some theoretical models (e.g. Hatano–Nelson model hatano1996localization and Su-Schrieffer-Heeger model su1979solutons ) in quantum physics.
@@
As non-Hermitian mechanical systems are also subjects of intensive study, some theories have been developed and validated. Melkani and Paulose studied time-modulated discrete mechanical systems based on space-time symmetry melkani2023space-time . Ghatak et al. discussed the non-Hermitian topology in an active mechanical metamaterial in analogy to the Su-Schrieffer-Heeger model ghatak2020observation . A mathematical theory for time-modulated discrete systems is also established by Ammari and Kosche ammari2024topological . Nassar et al. showed that mass-spring chains with space-time modulation break reciprocity nassar2018quantization . Salerno et al. studied complex frequency spectra of time-modulated mass-spring lattices and showed that they are associated with parametric instability of mechanical vibration salerno2016floquet . Kruss and Paulose discussed non-reciprocal one-way amplification of mechanical waves in time-modulated mass-spring chains using the complex Floquet analysis kruss2022nondispersive . Nevertheless, it remains unclear how the spectrum of time-modulated dynamic mechanical systems with OBCs is associated with that of periodic lattices.
@@
In this study, we propose a non-Bloch band theory for time-modulated discrete mechanical systems. Time-modulated mass-spring chains are formulated using the temporal Floquet theory. Our model is similar to the one analyzed in nassar2018quantization ; however, we generalize their Bloch–Floquet theory to complex frequency spectra for the analysis of the GBZ and non-Hermitian skin effects. We present numerical evidence that the conventional Bloch band theory is not valid for the time-modulated systems. Following yokomizo2019non-bloch , we define a GBZ and establish a non-Bloch band theory. This allows us to estimate eigenvalues of long chains in the complex frequency plane even in the presence of non-Hermitian skin modes that arise from the non-reciprocity.
@@ 2.1 Model
with constants $G_{j}>0$, $\gamma_{j}\in[0,G_{j})$, and $\phi_{j}\in[-\pi,\pi]$. Since all the coefficients $\mu_{j}$ are $T$-periodic, i.e., $\mu_{j}(t+T)=\mu_{j}(t)$ with $T=2\pi/\Omega$ for all $j=1,\ldots,M$ and $t\in\mathbb{R}$, we use the Floquet theory ammari2022nonreciprocal , ammari2023transmission , ammari2024topological and seek solutions $u^{j}$ of the form
@@
with unknown coefficients $U^{j}_{n},V^{j}_{n}\in\mathbb{C}$ ($n\in\mathbb{Z}$) for each $j=1,\ldots,M$. The coefficients should satisfy $(\omega+n\Omega)U^{j}_{n}=\mathrm{i}V^{j}_{n}$ for each $j$ and $n$. Note that the response $u^{j}$ is not in general time-harmonic with angular frequency $\omega$ despite the harmonic excitation $f_{j}$. In view of this, the factor $\omega$ in 6 is called quasifrequency ammari2023transmission .
@@
$F=$ $(0,\ldots,F_{1},\ldots,0,\ldots,0,\ldots,{F_{M}},\ldots,0)^{T}.$ (19)
@@
is not necessarily Hermitian despite that each matrix is Hermitian. Thus, we cannot always expect that eigenvalues $\omega$ of $\mathcal{H}$ are all real. However, due to the Hermiticity of the two matrices, it can be proved that the eigenvalues $\omega$ are real or come in complex-conjugate pairs $(\omega,\bar{\omega})$ tisseur2001quadratic .
@@ 2.2 Numerical example
Figure 2 (b) and (c) show the trajectories of eigenvalues $\omega$ of $\mathcal{H}$ as the modulation amplitude $\delta$ varies from 0 to $0.99$. As the spectrum is $\Omega$-periodic in the complex $\omega$-plane ammari2023transmission , we plot eigenvalues only within the range of $0\leq\mathrm{Re}[\omega]\leq\Omega$. The results indicate that all the eigenvalues are purely real when the modulation amplitude $\delta$ is less than $0.843$. Above this threshold, the eigenvalues come in a conjugate pair $(\omega,\bar{\omega})$ with nonzero imaginary part. The eigenvalues with negative and positive imaginary part are associated with modes decaying and growing in time, respectively (recall that the convention $\mathrm{e}^{-\mathrm{i}\omega t}$ is used). This means that the system undergoes a phase transition from stable to unstable states. This is a typical phenomenon in PT symmetric systems and called real-to-complex eigenvalue transition, pseudo-Hermiticity breaking, or PT phase transition mostafazadeh2002pseudo-hermiticity , melkani2023space-time .
@@
Figure 2: Verification of the Floquet theory for a time-modulated system. (a) Two mass-spring pairs. (b), (c) Trajectory of eigenvalues $\omega$ as the modulation amplitude $\delta$ varies from 0 to $0.99$. (d) Displacements $u^{1}$ and $u^{2}$ as a function of time $t$ when the initial conditions $u^{1}(0)=u_{0}$ and $u^{2}(0)=\partial_{t}u^{1}(0)=\partial_{t}u^{2}(0)=0$ are given.
Figure 3: Spectrum of $\mathcal{H}$ for a long-chain of mass-spring pairs. (a) A long chain consists of spatially periodic units. Its unit comprises three mass-spring pairs. (b) Time modulation of $\mu_{1}$, $\mu_{2}$, and $\mu_{3}$. (c) Distribution of eigenvalues $\omega$ for the system (a). Some representatives are denoted by the symbols A, B, C, D, and E with their conjugate pairs. (d) Eigenmodes corresponding to the representatives.
@@ 3 Non-Bloch band theory
In the previous section, we observed that the long chain with repeating units exhibit some special spectral structures and its eigenmodes can be localized at the edges. To analyze these unique properties, we approximate the long chain as a purely periodic system. The interesting point is that the standard Bloch band theory breaks down for our time-modulated system due to the non-Hermiticity of $\mathcal{H}$. In this section, we introduce the non-Bloch band theory kunst2019non-hermitian , yokomizo2019non-bloch with a GBZ yao2018edge , yang2020non-hermitian to study the spectrum of $\mathcal{H}$ for long chains with OBCs.
@@ 3.1 Bloch band theory with Brillouin zone
which depends on $k$. Let $\sigma(k)=\{\omega\in\mathbb{C}:\mathrm{det}(\omega I-\mathcal{H}_{k})=0\}$ be the spectrum of $\mathcal{H}_{k}$. The union $\sigma_{\mathrm{BZ}}:=\cup_{k\in\mathbb{R}}\ \sigma(k)$ is called a Bloch band. For the computation of $\sigma_{\mathrm{BZ}}$, it suffices to consider $k\in[-\pi,\pi]$ as the $2\pi$ shift does not affect the Bloch condition 24, i.e., $\mathcal{H}_{k+2\pi}=\mathcal{H}_{k}$ and $\sigma_{\mathrm{BZ}}=\cup_{k\in[-\pi,\pi]}\ \sigma(k)$. The range $\pi,\pi]$ is called the (first) Brillouin zone (BZ).
@@
However, this well-known fact does not always hold true for non-Hermitian systems zhang2022review . The Bloch condition 24 with real $k$ implicitly enforces solutions not to grow or decay in space, which is not a reasonable assumption from the results in Figure 3 (d).
To see this, we present a numerical example of the Bloch band $\sigma_{\mathrm{BZ}}$ in Figure 4. As shown in Figure 4 (a) and (b), we use the same modulation frequency $\Omega$ and material parameters $m_{j}$, $G_{j}$, $\gamma_{j}$, and $\phi_{j}$ in 21, 22 and 23. Figure 4 (c) and (d) show the Bloch band diagram on the BZ $\pi\leq k\leq\pi$. As the periodic system is also non-Hermitian, the eigenvalues of $\mathcal{H}_{k}$ are complex-valued in general. Notably, the bands are asymmetric with respect to the inversion $k\to-k$, which implies that the system is non-reciprocal. Figure 4 (e) shows the comparison between the bands $\sigma_{\mathrm{BZ}}$ and spectrum of the long chain $\sigma_{\mathrm{OBC}}$ in Figure 3 (c). From the result, we observe the obvious discrepancy between the two spectra. This implies that the Bloch band theory with the standard BZ does not work for our time-modulated system.
@@ 3.2 Non-Bloch band theory with generalized Brillouin zone
The system 28 yields a (generalized) linear eigenvalue problem in terms of $\omega$ for fixed $\beta\in\mathbb{C}$. When $\omega\in\mathbb{C}$ is given, this is a quadratic eigenvalue problem in terms of $\beta$, which can be recast into a linear eigenvalue problem tisseur2001quadratic .
For each $\omega\in\mathbb{C}$, let $\beta_{1}(\omega),\beta_{2}(\omega),\ldots,\beta_{2N}(\omega)$ be eigenvalues of 28, where $N:=6(2P+1)$ is the size of the whole matrix in 28. The eigenvalues are sorted such that $|\beta_{1}(\omega)|\leq|\beta_{2}(\omega)|\leq\ldots\leq|\beta_{2N}(\omega)|$. A generalized Brillouin zone (GBZ) yao2018edge , yang2020non-hermitian is defined by
@@
To see this, let $(\tilde{U}^{1}_{l},\tilde{U}^{2}_{l},\tilde{U}^{3}_{l},\tilde{V}^{1}_{l},\tilde{V}^{2}_{l},\tilde{V}^{3}_{l})$ be an eigenvector of 28 corresponding to $\beta_{l}$ for each $l=1,\ldots,2N$ (here we omit the dependence on $\omega$). Let us consider the finite-length system 12 with $F=0$, $M=3L$, and repeating constants, e.g., $m_{1}=m_{4}=\ldots=m_{3L-2}$, $m_{2}=m_{5}=\ldots=m_{3L-1}$, and $m_{3}=m_{6}=\ldots=m_{3L}$, where $L$ is the number of units. We assume that a solution of 12 can be written as
$\begin{pmatrix}U^{3j-2}\\ U^{3j-1}\\ U^{3j}\\ V^{3j-2}\\ V^{3j-1}\\ V^{3j}\end{pmatrix}=\sum_{l=1}^{2N}c_{l}\beta_{l}^{j-1}\begin{pmatrix}\tilde{U}^{1}_{l}\\ \tilde{U}^{2}_{l}\\ \tilde{U}^{3}_{l}\\ \tilde{V}^{1}_{l}\\ \tilde{V}^{2}_{l}\\ \tilde{V}^{3}_{l}\end{pmatrix}$ (33)
for all $j=1,\ldots,L$. The coefficients $c_{1},\ldots,c_{2N}\in\mathbb{C}$ are determined from the equations of motion and OBCs. In what follows, we limit ourselves to the homogeneous Dirichlet condition $(U^{1},V^{1})=(U^{3L},V^{3L})=(0,0)$. Other linear and homogeneous OBCs can be treated analogously.
From the OBC at the left end, we have
$\begin{pmatrix}0\\ 0\end{pmatrix}=\begin{pmatrix}U^{1}\\ V^{1}\end{pmatrix}=\sum_{l=1}^{2N}c_{l}\begin{pmatrix}\tilde{U}_{l}^{1}\\ \tilde{V}_{l}^{1}\end{pmatrix}$, (34)
or equivalently
$\begin{bmatrix}\varphi^{1}_{1}&\varphi^{1}_{2}&\cdots&\varphi^{1}_{2N}\end{bmatrix}\begin{pmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{2N}\end{pmatrix}=0$ (35)
with $\varphi^{1}_{l}:=(\tilde{U}_{l}^{1},\tilde{V}_{l}^{1})\in\mathbb{C}^{N/3}$ for each $l=1,\ldots,2N$. Analogously, the OBC at the right end yields
$\begin{bmatrix}\beta_{1}^{L-1}\psi^{1}_{1}&\beta_{2}^{L-1}\psi^{1}_{2}&\cdots&\beta_{2N}^{L-1}\psi^{1}_{2N}\end{bmatrix}\begin{pmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{2N}\end{pmatrix}=0,$ (36)
where $\psi^{1}_{1},\ldots,\psi^{1}_{2N}\in\mathbb{C}^{N/3}$ are defined by $\psi^{1}_{l}:=(\tilde{U}_{l}^{3},\tilde{V}_{l}^{3})$.
From the equation of motion for the leftmost mass (the first row of 12), we see that $(U^{1},V^{1})$ and $(U^{2},V^{2})$ satisfy the following condition:
$\begin{bmatrix}K_{1}&-\mathrm{i}m_{1}(\Omega B+\omega I)\\ \mathrm{i}m_{1}(\Omega B+\omega I)&m_{1}I\end{bmatrix}\begin{pmatrix}U^{1}\\ V^{1}\end{pmatrix}+\begin{bmatrix}-K_{1}&O\\ O&O\end{bmatrix}\begin{pmatrix}U^{2}\\ V^{2}\end{pmatrix}=\begin{pmatrix}0\\ 0\end{pmatrix}$, (37)
which can be rewritten as
$\begin{bmatrix}\varphi^{2}_{1}&\varphi^{2}_{2}&\cdots&\varphi^{2}_{2N}\end{bmatrix}\begin{pmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{2N}\end{pmatrix}=0,$ (38)
where $\varphi^{2}_{1},\ldots,\varphi^{2}_{2N}\in\mathbb{C}^{N/3}$ are defined by
$\varphi^{2}_{l}:=\begin{bmatrix}K_{1}&-\mathrm{i}m_{1}(\Omega B+\omega I)\\ \mathrm{i}m_{1}(\Omega B+\omega I)&m_{1}I\end{bmatrix}\begin{pmatrix}\tilde{U}_{l}^{1}\\ \tilde{V}_{l}^{1}\end{pmatrix}+\begin{bmatrix}-K_{1}&O\\ O&O\end{bmatrix}\begin{pmatrix}\tilde{U}_{l}^{2}\\ \tilde{V}_{l}^{2}\end{pmatrix}$ (39)
for all $l=1,\ldots,2N$.
The same argument can be applied to the second and last two masses. We finally arrive at the following $2N\times 2N$ linear system:
$\begin{bmatrix}\varphi^{1}_{1}&\varphi^{1}_{2}&\cdots&\varphi^{1}_{2N}\\ \varphi^{2}_{1}&\varphi^{2}_{2}&\cdots&\varphi^{2}_{2N}\\ \varphi^{3}_{1}&\varphi^{3}_{2}&\cdots&\varphi^{3}_{2N}\\ \beta_{1}^{L-1}\psi^{1}_{1}&\beta_{2}^{L-1}\psi^{1}_{2}&\cdots&\beta_{2N}^{L-1}\psi^{1}_{2N}\\ \beta_{1}^{L-1}\psi^{2}_{1}&\beta_{2}^{L-1}\psi^{2}_{2}&\cdots&\beta_{2N}^{L-1}\psi^{2}_{2N}\\ \beta_{1}^{L-1}\psi^{3}_{1}&\beta_{2}^{L-1}\psi^{3}_{2}&\cdots&\beta_{2N}^{L-1}\psi^{3}_{2N}\end{bmatrix}\begin{pmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{2N}\end{pmatrix}=\begin{pmatrix}0\\ 0\\ \vdots\\ 0\end{pmatrix}$, (40)
with some vectors $\varphi^{1}_{l}$, $\varphi^{2}_{l}$, $\varphi^{3}_{l}$, $\psi^{1}_{l}$, $\psi^{2}_{l}$, and $\psi^{3}_{l}$ ($l=1,\ldots,2N$), which are independent of $L$. The linear system 12 with $F=0$ admits a nontrivial solution when the determinant of the $2N\times 2N$ matrix in 40 vanishes. This is an algebraic equation in terms of $\beta_{1},\ldots,\beta_{2N}$. By the argument in the Supplemental Material of yokomizo2019non-bloch , the two values $\beta_{N}$ and $\beta_{N+1}$ must have the same modulus in order that they form a continuum band as $L\to\infty$.
The spectrum $\sigma_{\mathrm{GBZ}}$ is numerically identified based on the following algorithm. For various $\omega\in\mathbb{C}$, we solve the eigenvalue problem 28 in terms of $\beta$ and sort the eigenvalues $\beta_{1}(\omega),\ldots,\beta_{2N}(\omega)$ according to their modulus. If the difference $||\beta_{N}(\omega)|-|\beta_{N+1}(\omega)||$ is sufficiently small, the angular quasifrequency $\omega\in\mathbb{C}$ is regarded as a member of $\sigma_{\mathrm{GBZ}}$.